A projectile is launched into space with a given initial velocity. How can calculate its maximum height? Acceleration is not constant (must use Newton's law of gravitation)

Mass is not given


1
Expert's answer
2021-06-02T09:36:13-0400
"\frac{1}{2}mv^2-\frac{GmM}{R}=-\frac{GmM}{R+h}\\
\frac{1}{2}v^2-\frac{GM}{R}=-\frac{GM}{R+h}"